OpenStudy (anonymous):

find the exat value of 4 sqrt27/ sqrt3

OpenStudy (jiteshmeghwal9):

\[4\sqrt{27} \over \sqrt{3}\]\[4\sqrt{3 \times 3 \times 3} \over \sqrt{3}\]\[4 \times 3=12\]\[4\sqrt{3 \times 3}\cancel{\sqrt{3}} \over \cancel{\sqrt{3}}\]

OpenStudy (jiteshmeghwal9):

gt it @KiArUnsYu ??

OpenStudy (anonymous):

yes. thank you

OpenStudy (jiteshmeghwal9):

yw :)